OpenStudy (anonymous):

Find both X and Y intercepts of the equation. 4x+y=5

OpenStudy (tcp2315):

x int: 4x + (0) = 5 x= 5/4 (5/4, 0) y int: 4(0) +y =5 y=5 (0,5)